40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 347     Entry time: Fri Apr 19 09:21:07 2019
Author: Philip 
Type: Optics 
Category:  
Subject: Cleaning of OMC optics 

ach[Joe, Phillip, Koji, Stephen]

Work from 17.04.2019

First contact cleaning of OMC optics
We cleaned the OMC optic with first contact. After a first cleaning run all mirrors except for two looked
fine. One had some first contact residuals on the left at center height and another had some particle sitting
near the center area. As the ionized nitrogen gun didn't help we applied another round of first contact which resolved
the two issues. Unfortutanely the second run of cleaning again left some residuals of first contact at the edges.
We were able to peal these off with tweezers.

Placement of Optics at the breadboard
We cleaned the contact surfaces for the bonds using optic wipes and pure isopropanol. The placement wen't well for 3 of the 5 optics (low number of newtonian rings).
One was recleaned and placed on the breadboard again which seemed fine. For the 5th no newtonian rings could be seen (either verry ood or bad) we planed on trying it in the current set-up. Mirrors used can be seen in attachment 3.

 

Attachment 1: IMG_7877.JPG  1.854 MB  Uploaded Fri Apr 19 10:28:23 2019  | Hide | Hide all
IMG_7877.JPG
Attachment 2: IMG_7883.JPG  1.654 MB  Uploaded Fri Apr 19 10:34:27 2019  | Hide | Hide all
IMG_7883.JPG
Attachment 3: IMG_7884.JPG  1.214 MB  Uploaded Fri Apr 19 10:35:04 2019  | Hide | Hide all
IMG_7884.JPG
ELOG V3.1.3-